# StressPrompt: Does Stress Impact Large Language Models and Human Performance Similarly?

### v1

@@ Abstract
(b) Performance comparison of Llama-3-8B-Instruct and Phi-3-mini-4k-instruct across different stress levels on various benchmarks.
Figure 1: Performance analysis of LLMs under varying stress levels. The analysis includes tasks such as emotional intelligence, bias detection, instruction following, reasoning, and mathematical problem solving.
@@ Introduction
As shown in Figure 1(a), LLMs exhibit optimal performance under moderate stress, with noticeable declines in performance at both low and high-stress levels. Additionally, Figure 1(b) provides a comparative analysis across different benchmarks, illustrating the varied effects of stress on multiple dimensions of LLM capabilities. Our study makes several key contributions:
@@ Conclusion
In this study, we constructed a dataset named StressPrompt to induce varying levels of stress in LLMs. Our analysis shows that stress significantly affects the internal states of LLMs, with deeper layers exhibiting higher sensitivity to stress levels. Moderate stress can enhance performance in tasks such as instruction following, reasoning, and emotional intelligence, while higher stress levels negatively impact bias detection. We developed a stress scanner that effectively measures the impact of stress on LLMs’ internal states, providing a tool to evaluate model robustness and resilience. These findings highlight the necessity of adjusting stress levels based on task requirements to optimize LLM performance. Identifying optimal stress levels can improve the resilience and adaptability of AI systems, ensuring reliable performance under pressure. Future research could explore other psychological phenomena and their effects on LLMs, further bridging the gap between human intelligence and artificial intelligence.
Appendix A Appendix
Experimental Details
We evaluated the performance of several instruction-tuned large language models (LLMs) under varying stress conditions. The models evaluated include Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3-70B-Instruct (AI@Meta 2024), Phi-3-mini-4k-Instruct (Abdin et al. 2024), Qwen2-72B-Instruct, Qwen2-7B-Instruct (Yang et al. 2024), and Mistral-7B-Instruct-v0.3 (Jiang et al. 2023). To ensure reproducibility, the generation temperature was set to 00 during evaluations. Each model was configured with its specific dialogue tokens to clearly define conversational roles and instructions. StressPrompts were introduced as system instructions, with all other settings kept consistent to ensure fair comparisons.
We considered a diverse set of benchmarks to evaluate different dimensions of LLM capabilities, including emotional intelligence, bias detection, instruction following, reasoning ability, and mathematical problem-solving. The datasets used are: IFEval (Zhou et al. 2023), which evaluates a model’s ability to follow explicit instructions; BBH (Big Bench Hard) (Suzgun et al. 2022), a set of $23$ challenging tasks from the BigBench dataset; MATH (Hendrycks et al. 2021b), comprising high-school level competition problems; GPQA (Graduate-Level Google-Proof Q&A Benchmark) (Rein et al. 2023), a dataset with challenging questions crafted by PhD-level experts; MuSR (Multistep Soft Reasoning) (Sprague et al. 2023), featuring complex algorithmically generated problems; MMLU-P (Massive Multitask Language Understanding - Professional) (Wang et al. 2024b), a refined version of the MMLU dataset; EQ-Bench (Paech 2023), designed to evaluate emotional intelligence in LLMs; MMLU (Hendrycks et al. 2021a), a test covering $57$ diverse tasks; TruthfulQA (Lin, Hilton, and Evans 2021), a benchmark for measuring the truthfulness of generated answers; and ToxiGen (Hartvigsen et al. 2022), a large-scale dataset for adversarial and implicit hate speech detection.
The first five datasets constitute the new generation of the Open LLM Leaderboard [^2] https://huggingface.co/spaces/open-llm-leaderboard/open˙llm˙leaderboard, while the latter datasets provide additional perspectives on emotional intelligence, bias detection, hallucinations, and other critical aspects of model performance.
We utilized the lm_eval (Gao et al. 2023) framework for all evaluations, maintaining default settings to ensure reproducibility and broad applicability. This framework facilitates consistent and standardized evaluations across different models and datasets. To intuitively compare the impact of different prompts on the performance of LLMs, we set two baselines: the default prompt ’you are a helpful assistant’ and the chain-of-thought (CoT) prompting ’let’s think step by step’. Each of the StressPrompts was applied individually, and the LLMs’ responses were recorded and evaluated based on the specified metrics.
All evaluations were performed on NVIDIA A100 GPUs, with a maximum batch size of $16$ and adaptive execution to optimize performance. This comprehensive setup allowed us to systematically assess the impact of stress levels on LLM performance and draw meaningful insights into their robustness and resilience under different conditions.
To classify the stress levels of the prompts in our StressPrompt dataset, we recruited human participants offline. A total of $20$ participants were recruited and compensated at a rate that meets or exceeds the minimum hourly wage in our region. All participants provided informed consent prior to participating in the study, being informed about the purpose of the study, the procedures involved, and their rights as participants, including the right to withdraw at any time without any penalty.
Participants were provided with a set of 100 prompts, each designed to reflect different levels of stress based on established psychological frameworks. All participants rated the stress induced by each of the 100 prompts on a scale from 1 (minimal stress) to 10 (maximum stress). To ensure consistency and reliability in the ratings, the final stress level for each prompt was determined by averaging the ratings across all participants. All data collected from participants was anonymized to protect their privacy, with no personally identifiable information stored or shared. The dataset will be provided as an appendix for transparency and reproducibility. The distribution of stress levels across the prompts is shown in Figure 4.
Stress Level Base CoT 1 2 3 4 5 6 7 8 9 10
Llama-3-8B-Instruct
Mistral-7B-Instruct-v0.3
MMLU 60.60 60.63 60.25 ±0.37 60.16 ±0.28 60.53 ±0.44 60.48 ±0.14 60.50 ±0.15 60.38 ±0.10 60.38 ±0.33 60.39 ±0.24 60.34 ±0.27 60.28 ±0.11
BBH 44.86 45.12 44.83 ±0.24 44.74 ±0.21 44.71 ±0.18 44.61 ±0.18 44.70 ±0.45 44.80 ±0.14 44.70 ±0.17 44.55 ±0.29 44.42 ±0.26 44.40 ±0.33
GPQA 28.36 29.67 28.69 ±0.52 28.55 ±0.32 29.00 ±0.45 28.75 ±0.32 28.62 ±0.74 29.31 ±0.31 28.85 ±0.61 29.11 ±0.41 28.94 ±0.42 28.90 ±0.46
IFEval 58.27 57.31 57.20 ±1.61 57.58 ±1.08 58.22 ±1.02 58.08 ±1.16 58.88 ±1.63 58.66 ±1.13 58.42 ±1.10 57.95 ±1.12 57.02 ±1.28 57.34 ±0.74
MATH 2.43 2.60 2.24 ±0.29 2.13 ±0.47 2.45 ±0.29 2.47 ±0.12 2.66 ±0.35 2.61 ±0.46 2.70 ±0.28 2.62 ±0.26 2.32 ±0.19 2.35 ±0.15
MMLU-P 28.34 27.79 27.72 ±0.77 27.68 ±0.64 28.03 ±0.29 28.14 ±0.43 27.87 ±0.32 27.60 ±0.92 27.99 ±0.62 27.64 ±0.49 27.86 ±0.91 27.37 ±0.44
MuSR 36.34 35.54 36.66 ±0.78 36.83 ±0.37 36.75 ±0.58 36.77 ±0.67 36.54 ±0.35 36.92 ±0.48 36.95 ±0.60 36.88 ±0.60 37.12 ±0.85 37.08 ±0.48
Phi-3-mini-4k-Instruct
Qwen2-7B-Instruct
MMLU 69.91 69.96 69.43 ±0.27 69.55 ±0.21 69.64 ±0.25 69.71 ±0.19 69.77 ±0.08 69.69 ±0.07 69.71 ±0.11 69.74 ±0.10 69.71 ±0.12 69.67 ±0.13
BBH 50.82 51.21 50.14 ±0.36 50.22 ±0.61 50.48 ±0.37 50.74 ±0.27 50.65 ±0.14 50.50 ±0.37 50.27 ±0.56 50.23 ±0.53 50.04 ±0.69 49.89 ±0.82
GPQA 30.97 31.29 31.14 ±0.54 31.14 ±0.37 31.15 ±0.79 31.32 ±0.26 31.26 ±0.46 30.88 ±0.81 31.66 ±0.49 31.39 ±0.53 31.04 ±0.47 31.27 ±0.32
IFEval 60.79 61.99 61.86 ±0.87 62.16 ±0.73 62.18 ±1.01 62.34 ±0.96 63.01 ±0.65 62.57 ±1.03 62.67 ±1.13 62.92 ±0.78 62.68 ±1.00 63.01 ±0.89
MATH 0.03 0.04 0.00 ±0.00 0.00 ±0.01 0.01 ±0.02 0.02 ±0.03 0.06 ±0.06 0.03 ±0.03 0.01 ±0.03 0.00 ±0.00 0.00 ±0.00 0.00 ±0.00
MMLU-P 34.74 36.33 35.55 ±0.87 35.77 ±1.05 35.60 ±1.10 35.48 ±1.09 35.82 ±0.99 35.45 ±1.04 35.62 ±0.93 35.67 ±0.60 35.59 ±0.95 35.90 ±0.66
MuSR 39.01 39.67 40.02 ±0.64 40.05 ±0.47 39.59 ±0.67 39.79 ±0.49 39.38 ±0.44 39.59 ±0.39 39.62 ±0.48 39.54 ±0.43 39.83 ±0.36 39.49 ±0.42
Llama-3-70B-Instruct
MMLU 68.48 68.18 38.32 ±17.36 34.55 ±14.29 45.76 ±16.73 66.89 ±10.38 76.36 ±0.64 73.48 ±1.42 72.80 ±4.62 74.92 ±1.07 74.11 ±2.16 71.15 ±5.13
EQ-Bench 82.34 82.21 82.27 ±0.95 82.29 ±0.85 82.32 ±0.38 82.02 ±0.51 82.11 ±0.24 81.60 ±0.73 81.87 ±0.54 81.90 ±0.42 81.70 ±0.46 81.91 ±0.55
ToxiGen 87.44 86.80 83.33 ±2.69 83.53 ±3.93 86.30 ±0.59 87.05 ±0.87 86.65 ±0.70 86.83 ±0.94 86.00 ±1.16 86.48 ±1.04 85.59 ±0.99 85.62 ±0.87
TruthfulQA 62.57 60.37 62.56 ±0.92 62.92 ±1.01 62.74 ±1.04 62.68 ±0.96 63.40 ±0.47 63.82 ±0.38 63.72 ±1.06 63.43 ±0.79 64.08 ±1.16 63.35 ±0.87
Qwen2-72B-Instruct
MMLU 81.01 80.78 80.79 ±0.18 80.93 ±0.12 80.98 ±0.19 80.95 ±0.12 81.19 ±0.13 81.24 ±0.18 81.12 ±0.17 81.13 ±0.14 81.01 ±0.13 81.13 ±0.14
EQ-Bench 81.75 81.36 82.37 ±0.68 82.43 ±0.62 82.36 ±0.66 82.19 ±0.54 82.25 ±0.67 81.60 ±0.44 81.76 ±0.51 81.80 ±0.36 81.83 ±0.65 81.40 ±0.18
ToxiGen 85.00 84.93 84.93 ±0.91 84.54 ±1.43 85.14 ±1.13 86.13 ±0.72 86.38 ±0.69 86.38 ±0.71 86.26 ±0.81 86.43 ±0.57 86.40 ±0.73 86.55 ±0.59
TruthfulQA 70.84 71.07 71.07 ±0.84 71.17 ±1.15 71.50 ±0.94 71.52 ±0.69 71.42 ±0.39 71.64 ±0.26 71.82 ±0.44 71.65 ±0.41 71.98 ±0.47 71.58 ±0.40
Table A1: Performance of various models across different stress levels for various tasks. Values are averaged over multiple prompts and expressed with their respective standard deviations.

### v2

@@ Abstract
(b) Illustration of the Yerkes-Dodson law: human performance varies with stress levels, peaking at moderate stress and declining under low or high stress. [^1] https://en.wikipedia.org/wiki/Yerkes–Dodson˙law
Figure 1: Comparison of stress-level performance between LLMs and humans.
@@ Introduction
As shown in Figure 1(a), LLMs exhibit optimal performance under moderate stress, with noticeable declines in performance at both low and high-stress levels. Additionally, Figure 7 provides a comparative analysis across different benchmarks, illustrating the varied effects of stress on multiple dimensions of LLM capabilities. Our study makes several key contributions:
@@
Analysis Under Varying Stress Levels
Figure 7: Performance comparison across different stress levels on various benchmarks.
@@ Conclusion
Our study demonstrates that LLMs exhibit performance patterns closely resembling those of humans under varying stress levels. By constructing the StressPrompt dataset, we found that LLMs can map human-like relationships between stress and task performance. Moderate stress enhances capabilities such as reasoning and instruction following, while excessive stress impairs tasks like bias detection. This mapping enables LLMs to emulate human strategies in problem-solving, adapting stress levels to optimize performance. These findings suggest that large models have captured and operationalized human-like stress-performance dynamics, paving the way for more resilient and adaptive AI systems.
Acknowledgments
This research was funded by the Central Government-Guided Local Special Fund within the Beijing Science and Technology Program (Grant No.Z241100001324005).